# Control LLM: Controlled Evolution for Intelligence Retention in LLM

### v1

@@ Abstract
Extensive experiments demonstrate the effectiveness of Control LLM in both CPT and CSFT. On Llama3.1-8B-Instruct, it achieves significant improvements in mathematical reasoning ($+14.4\%$ on Math-Hard) and coding performance ($+10\%$ on MBPP-PLUS). On Llama3.1-8B, it enhances multilingual capabilities ($+10.6\%$ on C-Eval, $+6.8\%$ on CMMLU, and $+30.2\%$ on CMMLU-0shot-CoT). It surpasses existing methods and achieves SOTA among open-source models tuned from the same base model, using substantially less data and compute. Crucially, these gains are realized while preserving strong original capabilities, with minimal degradation ($<4.3\%\text{on MMLU}$) compared to $>35\%$ in open-source Math and Coding models. This approach has been successfully deployed in LinkedIn’s GenAI-powered job seeker and Ads unit products.
@@ 2 Related Work
Architecture-Based Methods: Adapter modules (Houlsby et al., 2019) insert task-specific layers, leaving the original model intact. LLaMA Pro (Wu et al., 2024) and SOLAR (Kim et al., 2023) expand model depth through additional transformer blocks or Depth Up-Scaling (DUS). Llama-MoE (Qu et al., 2024) partitions FFN or Attention layers into sparse experts and trains top-k gate.
@@
Control LLM: Combining architecture-based approaches with parameter isolation, Control LLM aligns hidden states of parallel pre-trained and expanded blocks via interpolation. Unlike many PEFT methods, it does not curtail learning capacity or demand original training data. As a result, it retains previously learned skills while accommodating new tasks.
@@ 3.2 Control LLM Architecture
An alignment mechanism fuses these branches’ outputs, integrating new representations with existing knowledge.
@@ 5 Discussion, Future Work, and Conclusion
Control LLM offers a novel solution to mitigate catastrophic forgetting (CF) in large language models by employing interpolation-based alignment between pre-trained and expanded representations. It achieves a robust balance between retaining prior knowledge and acquiring new skills, consistently outperforming baseline methods. The approach stabilizes fine-tuning, enabling smooth integration of new knowledge without degrading existing performance, while being robust to hyper-parameter sensitivity and avoiding loss spikes during training (Appendix A).
As a practical alternative to Mixture-of-Experts (MoE) architectures, Control LLM leverages a continuous and learnable interpolation mechanism, eliminating hard gating and expert silos. This design enables effective hidden-state alignment and seamless knowledge integration, making it particularly well-suited for dynamic adaptation to evolving data distributions.
Future Directions:
  • Design efficient architectures for seamless model merging.
  • Extend Control LLM to multimodal and reinforcement fine-tuning tasks.
  • Explore theoretical foundations of hidden-state alignment and its role in generalization.
  • Investigate advanced divergence loss formulations and adaptive weighting strategies.
In summary, Control LLM enhances new-task performance while preserving prior capabilities, providing a robust method for continuous learning in LLMs.

### v2

@@ Abstract
Extensive experiments demonstrate the effectiveness of Control LLM in both CPT and CSFT. On Llama3.1-8B-Instruct, it achieves significant improvements in mathematical reasoning ($+14.4\%$ on Math-Hard) and coding performance ($+10\%$ on MBPP-PLUS). On Llama3.1-8B, it enhances multilingual capabilities ($+10.6\%$ on C-Eval, $+6.8\%$ on CMMLU, and $+30.2\%$ on CMMLU-0shot-CoT). It surpasses existing methods and achieves SOTA among open-source models tuned from the same base model, using substantially less data and compute. Crucially, these gains are realized while preserving strong original capabilities, with minimal degradation ($<4.3\%\text{on MMLU}$) compared to $>35\%$ in open-source Math and Coding models. Control LLM has been successfully deployed in LinkedIn’s GenAI-powered job seeker and Ads unit products.
@@ 2 Related Work
Architecture-Based Methods: Adapter modules (Houlsby et al., 2019) insert task-specific layers, leaving the original model intact. LLaMA Pro (Wu et al., 2024) and SOLAR (Kim et al., 2023) expand model depth through additional transformer blocks or Depth Up-Scaling (DUS). Llama-MoE v2 (Qu et al., 2024) does the model upcycling by partitioning FFN or Attention layers into sparse experts and trains top-k gate.
@@
Control LLM combines an architecture-based model upcycling approach with parameter isolation. In addition, it aligns hidden states of parallel pre-trained and expanded blocks via interpolation and divergence loss, to the best of our knowledge, no prior work has explored. Unlike many PEFT methods, it does not curtail learning capacity or demand original training data. As a result, it retains previously learned skills while accommodating new tasks.
@@ 3.2 Control LLM Architecture
Interpolation with an alignment mechanism fuses these branches’ outputs, integrating new representations with existing knowledge, with no initial training performance decrease entailed by changing a trained network’s structure.
@@ 5 Discussion, Future Work, and Conclusion
Control LLM presents a more effective model upcycling architecture compared to MoE and other baselines. It leverages a continuous and learnable interpolation mechanism, eliminating hard gating and expert silos, enables effective hidden-state alignment and seamless knowledge integration, making it particularly well-suited for dynamic adaptation to evolving data distributions. However, it comes with the trade-off of increasing the number of activated parameters.
Future work includes:
  • Evolve the architecture to enable effective weight fusion.
  • Extend to multi-modality and RLHF.
  • Explore theoretical foundations of hidden-state alignment.
In summary, Control LLM offers a novel solution to mitigate catastrophic forgetting (CF) in large language models by employing interpolation-based alignment between pre-trained and expanded representations. It achieves a robust balance between retaining prior knowledge and acquiring new skills, consistently outperforming baseline methods with much less data and computation. The approach stabilizes fine-tuning, enabling smooth integration of new knowledge without degrading existing performance, while being robust to hyper-parameter sensitivity and avoiding loss spikes during training (Appendix A).